# Tomur: Traffic-Aware Performance Prediction of On-NIC Network Functions with Multi-Resource Contention

### v3

@@ 7 Evaluation
    Accuracy: Tomur achieves an average prediction error of 3.7% end-to-end, with 78.8% error reduction compared to state-of-the-art across NFs (7.2). Microbenchmarks further show that Tomur’s design choices on multi-resource contentions and changing traffic attributes are effective (7.3, 7.4).
@@
    Usecases: To see how Tomur can be beneficial in practice, we show two concrete usecases where it (1) facilitates resource-efficient placement decisions in NF scheduling, reducing NF SLA violations by 88.5% and 92.2% compared to the classical greedy-based approaches and SLOMO, and (2) enables fast performance diagnosis for NFs under contentions with 100% accuracy (7.5).
@@
    Overhead: Tomur’s offline adaptive profiling reduces profiling cost while maintaining high model accuracy, and takes less than 1 hour for each NF (7.6).
7.1 Methodology
We employ NVIDIA BlueField-2 (BF-2) to evaluate Tomur. A BF-2 SmartNIC has 8 ARMv8 A72 cores at 2.5GHz, 6MB L3 cache, 16GB DDR4 DRAM, dual ConnectX-6 100GbE ports, and hardware accelerators for regex and compression. The NF traffic is generated from a client machine with an AMD EPYC-7542 CPU with 32 cores at 2.9GHz and a ConnectX-6 100GbE NIC. Both the BF-2 server and client machine are connected to a Mellanox SN2700 switch. Simultaneous multi-threading and turbo boost are disabled.
Traffic profiles. We employ DPDK-Pktgen [23] to create various traffic profiles with different attributes, i.e., number of flows and packet sizes. The flow sizes follow a uniform distribution. In addition, we generate packet payloads using exrex [15] with diverse MTBR of conducting regular expression matches, e.g., 600 matches/MB, for NFs using regex accelerator. The rule sets are from [8].
Training. We generate various traffic profiles as said above and collect NF’s performance to prepare training data. The default profile has 16K flows, 1500B packet size, and the MTBR at 600 matches/MB. To account for multi-resource contention, we rely on the three synthetic benchmark NFs in 6 to generate training data in different contention levels. Unless otherwise stated, we train three models for each NF using different random seeds, and all prediction results are averaged by all three models for numerical stability.
Baseline. SLOMO [42] serves as our baseline. For each NF, we train a model using SLOMO’s gradient boost regression under the default traffic profile (Table 13, Appendix D). If in testing the traffic profile deviates from default, we employ SLOMO’s sensitivity extrapolation [^5] Sensitivity” refers to NF performance as a function of contention level for simplicity [28, 42, 44, 43]. to adapt the model (Section 6 in [42]). We validate that our models achieve the same level of prediction error as in [42] (Appendix A).
Metrics. We use the mean absolute percentage error (MAPE) commonly used in existing work [42, 25, 44] as the main metric of prediction accuracy. The ground truth is the NF’s real average throughput measured in identical deployment setup. To avoid the impact of test set size variations, we also report the number of prediction results experiencing errors within ±$\pm 5\%$ and ±$\pm 10\%$, and marked its percentage by the test set size as ±$\pm 5\%$ Acc. and ±$\pm 10\%$ Acc., respectively [59]. A larger value indicates higher accuracy.
NF. NF implementation details are given in 6. Each NF is allocated two dedicated CPU cores. Co-located NFs share the memory subsystem and hardware accelerators [41, 56, 60, 5, 4]. We always set traffic arrival rates (packets per second) high enough for NFs to achieve their maximum throughput.
NF SLOMO Tomur
@@ 7.1 Methodology
Table 2: Prediction accuracy comparison under both multi-resource contention and varying traffic attributes. On average Tomur reduces MAPE by 78.8% compared to SLOMO, at 3.7% and 17.5%, respectively. Unless otherwise stated, rows of table are sorted in ascending order of Tomur’s MAPE.
7.2 Overall Accuracy
We first evaluate the overall prediction accuracy of Tomur under multi-resource contention and varying traffic attributes. Each target NF is co-located with up to three other NFs and we numerate all possible combinations of NFs. We also apply 9 distinct traffic profiles for each NF. The results are aggregated under all traffic profiles, as shown in Table 2. We can see that Tomur averagely exhibits 3.7% MAPE, 70.8% ±$\pm 5\%$ Acc. and 95.9% ±$\pm 10\%$ Acc., compared to SLOMO’s 17.5% MAPE, 50.0% ±$\pm 5\%$ Acc. and 72.7% ±$\pm 10\%$ Acc., demonstrating Tomur’s superior prediction accuracy. Specifically, Tomur has the most significant gains for IPTunnel, FlowMonitor, FlowStats, and NIDS that use multiple resources and/or is sensitive to traffic attributes. Meanwhile, Tomur achieves the same accuracy as SLOMO for ACL because it is very lightweight and insensitive to traffic attributes. Note that the high prediction accuracy for such NFs is aligned with that from previous studies[42, 28, 44].
Next, we microbenchmark Tomur’s major design choices to better understand the benefits they each bring.
7.3 Deep-Dive: Multi-Resource Contention
First we look at how Tomur’s per-resource modeling and composition approach handles multi-resource contention. We fix the traffic profile to be the default in this section to isolate its impact. We choose FlowMonitor and NIDS that utilize multiple resources, and co-run each with mem-bench and regex-bench with varying contention levels. Table 3 presents the comparison. We can observe that Tomur reduces the MAPE by 44.2% and 17.1% for FlowMonitor and NIDS, respectively, To better analyze the source of accuracy gains, we zoom in on FlowMonitor and vary the regex contention level it receives in two ranges: low range (MTBR $\leq$ 600 matches/MB), and high range (MTBR $>$ 600 matches/MB). Figure 7(a) depicts the distribution of the absolute percentage errors. Tomur maintains low errors as contention rises with median errors consistently below 6.0%. With low contention, SLOMO also achieves high accuracy with median error at 2.5%, because in this case multi-resource contention effectively reduces to single-resource (memory for FlowMonitor) contention. Here Tomur’s error is actually slightly higher that SLOMO, because SLOMO enjoys the same amount of training data with Tomur but concentrate on one fixed traffic profile. During high contention, SLOMO’s inability to model contention on these two resources simultaneously lead to high median errors at 24.4%.
NF SLOMO Tomur
@@ 7.3 Deep-Dive: Multi-Resource Contention
Table 3: Prediction accuracy of SLOMO and Tomur when the NF runs under only multi-resource contention. Traffic profile is fixed to the default one.
Figure 7: The distribution (box and whisker plot) of the absolute percentage errors of (a) multi-resource contention with different regex contention levels, and (b) memory-only contention with different ranges of variations in the flow count. We show the value of median error in both figures.
To further evaluate the use of Tomur’s composition design based on execution pattern (4.2), we write two synthetic NFs, NF1 and NF2, with NF1 using memory and regex, while NF2 adds the compression accelerator. Each has both a pipeline and run-to-completion version. We compare Tomur to the simple sum and min composition (recall 2.2), where they use the same per-resource models trained specifically for NF1 and NF2. As shown in Table 4, Tomur attains the best accuracy across all cases and resource usage patterns, with MAPE lower than 2%. This gain essentially comes from properly modeling NF’s execution pattern to capture the end-to-end impact of multi-resource contention. Tomur is on par with min composition for pipeline NFs, which is consistent with Equation (7).
@@
Table 4: Prediction error of different multi-resource composition approaches for different execution patterns.
NF SLOMO Tomur
@@
Table 5: Prediction accuracy of SLOMO and Tomur when target NF runs under memory-only contention and dynamic traffic profiles.
7.4 Deep-Dive: Traffic Attributes
We now move to examine Tomur’s design on traffic-aware modeling. We choose traffic-sensitive NFs and co-run each with mem-bench on a single BF-2. Since SLOMO only considers memory contention, we set a fixed contention level in memory, exclude other resource contention, and generate 100 traffic profiles by randomly changing number of flows, packet size, and MTBR when applicable for each NF. Table 5 shows the results aggregated from all profiles. Again Tomur attains superior accuracy over SLOMO across the board, with over 90% in ±$\pm 10\%$ Acc. and $<$5% MAPE for most NFs.
We then zoom into one particular attribute, flow count, which SLOMO also specifically models. We vary the flow count between training and testing across two ranges: low range where it changes by at most 20%, and high range ($>$20%). Figure 7(b) displays the distribution of absolute percentage error in this case. Tomur maintains low errors consistently, with a median at 4.7%, highlighting the benefits of our traffic-aware modeling (5). Under low-range variations, SLOMO exhibits high accuracy with sensitivity extrapolation [42]. However, as traffic profiles undergo more significant changes, SLOMO suffers from high prediction errors, with 13.4% at the median. This is consistent with [42]: the extrapolation only works when the NF’s sensitivity profile in training has enough overlap with that under testing traffic [42], which corresponds to low range profiles here.
7.5 Tomur Use Cases
We now illustrate Tomur’s practical benefits through two use cases: (1) It enables contention-aware scheduling of NFs to improve resource utilization; (2) it facilitates performance diagnosis for NFs with dynamic traffic.
7.5.1 Contention-Aware Scheduling
We consider the scenario in which the operator places the NFs as they arrive to a group of SmartNICs to maximize resource utilization (i.e., minimize SmartNICs used) while maintaining their SLAs. SLA here is defined as the maximum allowed throughput drop relative to the baseline when the NF runs solo. Given that the offline version of this problem is NP-complete bin-packing [58, 27], we follow previous works[42, 56, 61] and consider online heuristics that deploy the NFs one by one. Specifically, we compare the following strategies: (1) Monopolization, which forbids co-location of NFs; (2) Greedy, which places an NF onto the SmartNIC with most available resources[41, 53]; (3) Contention-aware, which first predicts the performance of all NFs on a SmartNIC if the current NF gets deployed onto this SmartNIC, and then deploys the NF onto the SmartNIC if no SLA violation is predicted. Both Tomur and SLOMO can provide such contention-aware predictions. A new SmartNIC is added to the cluster when there is no feasible placement.
Table 6 compares the above strategies over 100 random sequences of 500 NF arrivals each. Each NF is assigned the default traffic profile, and its SLA is set to 5-20% throughput drop. We examine resource wastage, i.e., how many additional NICs are used against the optimal plan found by exhaustive search, and the corresponding SLA violations. We observe that Tomur minimizes resource wastage to merely 0.5% and reduces SLA violations by 88.5% and 92.2% over Greedy and SLOMO on average. The near-optimal performance of Tomur illustrates its potential in coordinating NF scheduling in a real-world scenario.
@@ 7.5.1 Contention-Aware Scheduling
Tomur 0.5 1.9
Table 6: Tomur’s usecase in contention-aware scheduling comparing to other baseline strategies. SLOMO’s negative resource overhead stems from erroneous placements compared to the optimal deployment.
7.5.2 Performance Diagnosis
In practice, performance diagnosis has important values as it allows programmers to systematically explore the design spaces, identify performance bottlenecks and optimization opportunities, and even provide early-stage insights/guidances on next-generation SmartNIC [45, 42]. Here we show another usecase of Tomur in diagnosing performance bottlenecks in NFs with dynamic traffic, when the bottleneck may shift across resources.
As Table 7 shows we deploy FlowStats, FlowMonitor, and IPComp Gateway that all use regex accelerator. We co-run each of them with mem-bench and regex-bench and adjust the MTBR from 0 to 1100 matches/MB while keeping memory contention levels unchanged, and manually analyze the actual its performance bottleneck using the hotspot analysis function of perf-tools [19]. This is the ground truth. We then calculate the percentage of correct identification of bottleneck using our prediction models. We can see Tomur accurately identifies bottleneck for all three NFs with its multi-resource performance modeling, while SLOMO only works for FlowStats, as it is always bottlenecked on memory. FlowMonitor and IPComp Gateway’s bottleneck actually shifts with traffic. For example, we observe that FlowMonitor’s bottleneck is memory with MTBR at 80 matches/MB, but changes to regex with MTBR at 1000 matches/MB (default traffic profile). Thus this usecase demonstrates Tomur’s capability in pinpointing NF bottlenecks with dynamic traffic.
@@ 7.5.2 Performance Diagnosis
Table 7: Percentage of correct identifications of performance bottleneck using SLOMO and Tomur.
@@
Table 8: Profiling cost and model accuracy using full, random and Tomur’s adaptive profiling. P.C. refers to profiling cost.
Figure 8: Prediction error on FlowClassfier using full, random and Tomur’s adaptive profiling. We change the profiling quota of random and adaptive profiling to 0.5$\times$ and 1.5$\times$ of that used in Table 8.
7.6 Offline Profiling
Lastly we examine the adaptive profiling design of Tomur.
Adaptive profiling. To evaluate Tomur’s adaptive profiling, we select traffic-sensitive NFs from Table 1, and train them using three different approaches — full profiling, random profiling, and adaptive profiling (5.2). For random profiling and Tomur’s adaptive profiling, we set the same number of training data points (a.k.a. profiling quota) to ensure a fair comparison. For full profiling, we use 80% of the profiled data for training, and the remaining 20% for testing. Profiling cost is represented by the number of training data samples normalized against adaptive/random profiling’s quota. We observe from Table 8 that Tomur’s adaptive profiling offers comparable accuracy to full profiling which uses 3200$\times$ more data. Compared to random, adaptive profiling significantly enhances accuracy within the same profiling quota. Notably, it achieves a maximum 35.5% MAPE reduction on FlowTracker and a 72% improvement in $\pm$10Acc. on FlowClassifier over random. We further show the benefit of adaptive profiling using FlowClassifier as an example in Figure 8. We adopt same setting as Table 8 but change the profiling quota to 0.5$\times$ and 1.5$\times$. Figure 8 reveals that by increasing current profiling quota by 50% (still $\sim$2100$\times$ less than the profiling cost of full profiling), adaptive profiling achieve similar error compared to full profiling, at 2.4% and 2.3%, respectively, while random profiling does not exhibit accuracy improvement since performance-significant ranges of traffic attribute is still not covered in the training data.
Time cost of profiling. As discussed in 6, Tomur requires offline profiling for each NF to build per-resource models and identify execution pattern. This process primarily involves collection of: (1) the contention level of synthetic benchmark NFs, and (2) contention level and sensitivity profiles of target NFs. Across our experiments, we find that on average 1.6 hours and 0.5 hour are spent on profiling the contention level of synthetic benchmark NFs (mem-bench, etc.) and the contention level and sensitivity profiles of an NF, respectively. These time investments are acceptable since profiling is a one-time effort.
8 Discussion
Can Tomur be generalized to other SmartNICs? Tomur can be generalized to other SoC SmartNICs due to the similar architecture of their hardware accelerators and memory subsystem. To quickly validate such generalizability, we collect data and train performance models of a Firewall NF [24] running on an AMD Pensando SmartNIC [13]. The NF conducts a flow walk on hardware flow table and updates entry metadata upon matching against flows in the input traffic, thus requires memory and regex accelerator. As shown in Table 9, the average prediction error of Tomur is 0.9%, which is 17.5% lower compared to that of SLOMO. Such preliminary result reflects that applying Tomur to other SoC SmartNICs is feasible. However, for SmartNICs whose architecture significantly deviates from SoC SmartNICs, e.g., on-path SmartNICs, we require more investigations on the contention behavior of their specialized hardware resources [38, 30]. We leave this as future work.
NF SLOMO Tomur

### v5

@@ 7. Evaluation
    Accuracy: Yala achieves an average prediction error of 3.7% end-to-end, with 78.8% error reduction compared to state-of-the-art across NFs (7.2). Microbenchmarks further show that Yala’s design choices on multi-resource contentions and changing traffic attributes are effective (7.3, 7.4).
@@
    Usecases: To see how Yala can be beneficial in practice, we show two concrete usecases where it (1) facilitates resource-efficient placement decisions in NF scheduling, reducing NF SLA violations by 88.5% and 92.2% compared to the classical greedy-based approaches and SLOMO, and (2) enables fast performance diagnosis for NFs under contention with 100% accuracy (7.5).
@@
    Overhead: Yala’s offline adaptive profiling reduces profiling cost while maintaining high model accuracy (7.6).
7.1. Methodology
We employ NVIDIA BlueField-2 (BF-2) to evaluate Yala. A BF-2 SmartNIC has 8 ARMv8 A72 cores at 2.5GHz, 6MB L3 cache, 16GB DDR4 DRAM, dual ConnectX-6 100GbE ports, and hardware accelerators for regex and compression. The NF traffic is generated from a client machine with an AMD EPYC-7542 CPU with 32 cores at 2.9GHz and a ConnectX-6 100GbE NIC. Both the BF-2 server and client machine are connected to a Mellanox SN2700 switch.
Traffic profiles. We employ DPDK-Pktgen (pktgen, ) to create various traffic profiles with different attributes, i.e., number of flows and packet sizes. In addition, we generate packet payloads using exrex (exrex, ) with diverse MTBR of conducting regular expression matches for NFs using regex accelerator. The rule sets are from (l7filter, ).
Baseline. SLOMO (slomo20, ) serves as our baseline. For each NF, we train a model using SLOMO’s gradient boost regression under the default traffic profile, which has 16K flows, 1500B packet size, and the MTBR at 600 matches/MB. If in testing the traffic profile deviates from default, we employ SLOMO’s sensitivity extrapolation [^5] Sensitivity” refers to NF performance as a function of contention level for simplicity (dobre12, ; slomo20, ; bubbleup11, ; opportunity11, ). to adapt the model (Section 6 in (slomo20, )). We validate that our models achieve the same level of prediction error as in (slomo20, ).
Metrics. We report mean absolute percentage error (MAPE) (slomo20, ; prophet17, ; bubbleup11, ) as the main metric of prediction accuracy. We also report ±$\pm 5\%$ Acc. and ±$\pm 10\%$ Acc. to avoid the impact of test set size variations (nnmeter, ).
7.2. Overall Accuracy
We first evaluate the overall prediction accuracy of Yala under multi-resource contention and varying traffic attributes. Each target NF is co-located with up to three other NFs and we numerate all possible combinations of NFs. We also apply 9 distinct traffic profiles for each NF. The results are aggregated under all traffic profiles, as shown in Table 2. We can see that Yala averagely exhibits 3.7% MAPE, 70.8% ±$\pm 5\%$ Acc. and 95.9% ±$\pm 10\%$ Acc., compared to SLOMO’s 17.5% MAPE, 50.0% ±$\pm 5\%$ Acc. and 72.7% ±$\pm 10\%$ Acc., demonstrating Yala’s superior prediction accuracy. Specifically, Yala has the most significant gains for IPTunnel, FlowMonitor, FlowStats, and NIDS that use multiple resources and/or is sensitive to traffic attributes. Meanwhile, Yala achieves the same accuracy as SLOMO for ACL because it is very lightweight and insensitive to traffic attributes. Note that the high prediction accuracy for such NFs is aligned with that from previous studies (slomo20, ; dobre12, ; bubbleup11, ).
Next, we microbenchmark Yala’s major design choices to better understand the benefits they each bring.
NF SLOMO Yala
@@ 7.2. Overall Accuracy
Table 2. Prediction accuracy comparison under both multi-resource contention and varying traffic attributes. On average Yala reduces MAPE by 78.8% compared to SLOMO, at 3.7% and 17.5%, respectively. Unless otherwise stated, rows of table are sorted in ascending order of Yala’s MAPE.
7.3. Deep-Dive: Multi-Resource Contention
First we look at how Yala’s per-resource modeling and composition approach handles multi-resource contention. We fix the traffic profile to be the default in this section to isolate its impact. We choose FlowMonitor and NIDS that utilize multiple resources, and co-run each with mem-bench and regex-bench with varying contention levels. Table 3 presents the comparison. We can observe that Yala reduces the MAPE by 44.2% and 17.1% for FlowMonitor and NIDS, respectively, To better analyze the source of accuracy gains, we zoom in on FlowMonitor and vary the regex contention level it receives in two ranges: low range (MTBR $\leq$ 600 matches/MB), and high range (MTBR $>$ 600 matches/MB). Figure 7(a) depicts the distribution of the absolute percentage errors. Yala maintains low errors as contention rises with median errors consistently below 6.0%. With low contention, SLOMO also achieves high accuracy with median error at 2.5%, because in this case multi-resource contention effectively reduces to single-resource (memory for FlowMonitor) contention. Here Yala’s error is actually slightly higher that SLOMO, because SLOMO enjoys the same amount of training data as Yala but concentrate on one fixed traffic profile. During high contention, SLOMO’s inability to model contention on these two resources simultaneously lead to high median errors at 24.4%.
NF SLOMO Yala
@@ 7.3. Deep-Dive: Multi-Resource Contention
Table 3. Prediction accuracy of SLOMO and Yala when the NF runs under only multi-resource contention. Traffic profile is fixed to the default one.
Figure 7. The distribution (box and whisker plot) of the absolute percentage errors of (a) multi-resource contention with different regex contention levels, and (b) memory-only contention with different ranges of variations in the flow count. We show the value of median error in both figures.
[…]
To further evaluate the use of Yala’s composition design based on execution pattern (4.2), we write two synthetic NFs, NF1 and NF2, with NF1 using memory and regex, while NF2 adds the compression accelerator. Each has both a pipeline and run-to-completion version. We compare Yala to the simple sum and min composition (recall 2.2.1), where they use the same per-resource models trained specifically for NF1 and NF2. As shown in Table 4, Yala attains the best accuracy across all cases and resource usage patterns, with MAPE lower than 2%. This gain essentially comes from properly modeling NF’s execution pattern to capture the end-to-end impact of multi-resource contention.
@@
Table 4. Prediction error of different multi-resource composition approaches for different execution patterns.
7.4. Deep-Dive: Traffic Attributes
We now move to examine Yala’s design on traffic-aware modeling. We choose traffic-sensitive NFs and co-run each with mem-bench on a single BF-2. Since SLOMO only considers memory contention, we set a fixed contention level in memory, exclude other resource contention, and generate 100 traffic profiles by randomly changing number of flows, packet size, and MTBR when applicable for each NF. Table 5 shows the results aggregated from all profiles. Again Yala attains superior accuracy over SLOMO across the board, with over 90% in ±$\pm 10\%$ Acc. and $<$5% MAPE for most NFs.
We then zoom into one particular attribute, flow count, which SLOMO also specifically models. We vary the flow count between training and testing across two ranges: low range where it changes by at most 20%, and high range ($>$20%). Figure 7(b) displays the distribution of absolute percentage error in this case. Yala maintains low errors consistently, with a median at 4.7%, highlighting the benefits of our traffic-aware modeling (5). Under low-range variations, SLOMO exhibits high accuracy with sensitivity extrapolation (slomo20, ). However, as traffic profiles undergo more significant changes, SLOMO suffers from high prediction errors, with 13.4% at the median. This is consistent with (slomo20, ): the extrapolation only works when the NF’s sensitivity profile in training has enough overlap with that under testing traffic (slomo20, ), which corresponds to low range profiles here.
NF SLOMO Yala
@@ 7.4. Deep-Dive: Traffic Attributes
Table 5. Prediction accuracy of SLOMO and Yala when target NF runs under memory-only contention and dynamic traffic profiles.
7.5. Yala Use Cases
We now illustrate Yala’s practical benefits through two use cases: (1) It enables contention-aware scheduling of NFs to improve resource utilization; (2) it facilitates performance diagnosis for NFs with dynamic traffic.
7.5.1. Contention-Aware Scheduling
We consider the scenario in which the operator places the NFs as they arrive to a group of SmartNICs to maximize resource utilization (i.e., minimize SmartNICs used) while maintaining their SLAs. SLA here is defined as the maximum allowed throughput drop relative to the baseline when the NF runs solo. Given that the offline version of this problem is NP-complete bin-packing (binpack, ; binpack1, ), we follow previous works (slomo20, ; resq18, ; octans19, ) and consider online heuristics that deploy the NFs one by one. Specifically, we compare the following strategies: (1) Monopolization, which forbids co-location of NFs; (2) Greedy, which places an NF onto the SmartNIC with most available resources (e319, ; meili23, ); (3) Contention-aware, which first predicts the performance of all NFs on a SmartNIC if the current NF gets deployed onto this SmartNIC, and then deploys the NF onto the SmartNIC if no SLA violation is predicted. Both Yala and SLOMO can provide such contention-aware predictions. A new SmartNIC is added to the cluster when there is no feasible placement.
Table 6 compares the above strategies over 100 random sequences of 500 NF arrivals each. Each NF is assigned the default traffic profile, and its SLA is set to 5-20% throughput drop. We examine resource wastage, i.e., how many additional NICs are used against the optimal plan found by exhaustive search, and the corresponding SLA violations. We observe that Yala minimizes resource wastage to merely 0.5% and reduces SLA violations by 88.5% and 92.2% over Greedy and SLOMO on average. The near-optimal performance of Yala illustrates its potential in coordinating NF scheduling in a real-world scenario.
@@ 7.5.1. Contention-Aware Scheduling
Yala 0.5 1.9
Table 6. Yala’s usecase in contention-aware scheduling compared to other baseline strategies. SLOMO’s negative resource overhead stems from erroneous placements compared to the optimal deployment.
@@
Table 7. Percentage of correct identifications of performance bottleneck using SLOMO and Yala.
7.5.2. Performance Diagnosis
In practice, performance diagnosis has important values as it allows programmers to systematically explore the design spaces, identify performance bottlenecks and optimization opportunities, and even provide early-stage insights/guidances on next-generation SmartNIC (morpheus22, ; slomo20, ). Here we show another usecase of Yala in diagnosing performance bottlenecks in NFs with dynamic traffic, when the bottleneck may shift across resources.
As Table 7 shows we deploy FlowStats, FlowMonitor, and IPComp Gateway that all use regex accelerator. We co-run each of them with mem-bench and regex-bench and adjust the MTBR from 0 to 1100 matches/MB while keeping memory contention levels unchanged, and manually analyze the actual its performance bottleneck using the hotspot analysis function of perf-tools (perf, ). This is the ground truth. We then calculate the percentage of correct identification of bottleneck using our prediction models. We can see Yala accurately identifies bottleneck for all three NFs with its multi-resource performance modeling, while SLOMO only works for FlowStats, as it is always bottlenecked on memory. FlowMonitor and IPComp Gateway’s bottleneck actually shifts with traffic. For example, we observe that FlowMonitor’s bottleneck is memory with MTBR at 80 matches/MB, but changes to regex with MTBR at 1000 matches/MB (default traffic profile). Thus this usecase demonstrates Yala’s capability in pinpointing NF bottlenecks with dynamic traffic.
7.6. Adaptive Profiling
Lastly we examine the adaptive profiling design of Yala. We select traffic-sensitive NFs from Table 1, and train them using three different approaches — full profiling, random profiling, and adaptive profiling (5.2). For random profiling and Yala’s adaptive profiling, we set the same number of training data points (a.k.a. profiling quota) to ensure a fair comparison. For full profiling, we use 80% of the profiled data for training, and the remaining 20% for testing. Profiling cost is represented by the number of training data samples normalized against adaptive/random profiling’s quota. We observe from Table 8 that Yala’s adaptive profiling offers comparable accuracy to full profiling which uses 3200$\times$ more data [^6] This is because we use 16 values for packet sizes and 200 values for the number of flows in full profiling. For each (packet size, number of flows) tuple, full profiling repeats profiling over a set of random memory contention levels.. Compared to random, adaptive profiling significantly enhances accuracy within the same profiling quota. We further show the benefit of adaptive profiling using FlowClassifier as an example in Figure 8. We adopt the same setting as Table 8 but change the profiling quota to 0.5$\times$ and 1.5$\times$. Figure 8 reveals that by increasing current profiling quota by 50% (still $\sim$2100$\times$ less than the profiling cost of full profiling), adaptive profiling achieve similar error compared to full profiling, at 2.4% and 2.3%, respectively, while random profiling does not exhibit accuracy improvement since performance-significant ranges of traffic attribute is still not covered in the training data.
@@ 7.6. Adaptive Profiling
Table 8. Profiling cost and model accuracy using full, random and Yala’s adaptive profiling. P.C. refers to profiling cost.
Figure 8. Prediction error on FlowClassfier using full, random and Yala’s adaptive profiling. We change the profiling quota of random and adaptive profiling to 0.5$\times$ and 1.5$\times$ of that used in Table 8.
[…]
Time cost of profiling. As discussed in 4, Yala requires offline profiling for each NF to build per-resource models and identify the execution pattern. This process primarily involves the collection of: (1) the contention level of synthetic benchmark NFs (mem-bench, etc.), and (2) the contention level and sensitivity profiles of the target NF. Across our experiments, we find that on average 1.6 hours and 0.5 hours for each, respectively. These time investments are acceptable since profiling is a one-time effort.
8. Discussion
We discuss a few issues one may have regarding Yala.
What if the configuration of an NF changes? It is possible that an NF’s configuration is adjusted for various reasons, e.g. adjusting ACL rules in a Firewall NF to apply new policies. Such changes can cause performance characteristics to change, making the existing model inaccurate. To make Yala “configuration-aware”, one may adopt a similar approach of traffic attributes, i.e., extracting “configuration attributes” for an NF and integrating it into the per-resource models. We leave this as future work.
Can Yala be generalized to other SmartNICs? Yala can be generalized to other SoC SmartNICs due to the similar architecture of their hardware accelerators and memory subsystem. To quickly validate such generalizability, we collect data and train performance models of a Firewall NF (db23, ) running on an AMD Pensando SmartNIC (pensando, ). The NF conducts a flow walk on hardware flow table and updates entry metadata upon matching against flows in the input traffic. As shown in Table 9, the average prediction error of Yala is 0.9%, which is 17.5% lower compared to that of SLOMO. This result reflects that applying Yala to other SoC SmartNICs is feasible. However, for SmartNICs like on-path SmartNICs whose architecture significantly deviates from SoC SmartNICs, more investigation is needed into the contention behavior of their specialized hardware resources (switchvm24, ; menshen22, ).
NF SLOMO Yala